VUV - EUV FOURIER TRANSFORM ABSORPTION SPECTROSCOPY WITHOUT A BEAM SPLITTER

D. JOYEUX, D. PHALIPPOU, J.-C. RODIER, Laboratoire Charles Fabry (Institut d’Optique et CNRS), Bât. 503, Université Paris-Sud, 91405 Orsay Cedex, France; N. DE OLIVEIRA, F. PO-LACK, Synchrotron Soleil, Orme des Merisiers, St Aubin BP 48, 91192 Gif sur Yvette Cedex, France; L. NAHON, LURE, Bât. 209 D, Université Paris-Sud, 91405 Orsay Cedex, France; \underline{M. VERVLOET}, Laboratoire de Photophysique Moléculaire du CNRS, Bât. 350, Université Paris-Sud, 91405 Orsay Cedex, France.

A new Fourier transform spectrometer without a beam splitter has been designed to operate on a VUV synchrotron beam line. It is based on a modified Fresnel bimirror configuration controlled by an original optical system. This system keeps the mirror tilt error to a negligible value (much lower than 1 μrad ) during the mirror translation. It also controls the sampling of the interferogram (1 sample per 15 nm) with the required accuracy, producing a sampling error less than a few nanometers over the scanning range. An absorption spectrum of the Schumann-Runge band system of molecular oxygen has been recorded at around 200 nm with a resolving power of about 150000. It shows great promise in terms of extension of the technique towards the VUV range. This spectrum will be presented and compared to the calculated spectrum obtained by using a published database^{a}.

ahttp://cfa-www.harvard.edu/amdata/ampdata/o2pub92/S-R.html